An oil drop of mass 10-12 gm is floating on the free surface of a liquid. At any instant the position of the drop can be determined within the error of 10-4 cm. what will be the error in measurement of its velocity.


1
Expert's answer
2021-03-29T09:01:37-0400

According to the Heisenberg's uncertainty principle


"\Delta p\Delta x \ge\dfrac{\hbar}{2}"

whereย "\hbar \approx 1.05\times 10^{-34}J\cdot s"ย is the reduced Planck constant, "\Delta x = 10^{-4}cm = 10^{-6}m" is the position error, and "\Delta p" is the momentum error. Since the mass of the drop is "m = 10^{-12}g = 10^{-15}kg", then


"\Delta p = m\Delta v"

where "\Delta v" is the error in measurement of drop's velocity. Thus, obtian:


"m\Delta v\Delta x \ge\dfrac{\hbar}{2} \\
\Delta v \ge\dfrac{\hbar}{2m\Delta x}\\
\Delta v \ge\dfrac{1.05\times 10^{-34}}{2\cdot 10^{-15}\cdot 10^{-6}} = 5.25\times 10^{-14}m/s"

Answer. "\ge5.25\times 10^{-14}m/s".